OpenStudy (anonymous):

convert prefix 4500 kgm^-3 ---> ? gcm^-3

OpenStudy (anonymous):

4500kg m^(-3)=4500*1000/10^6=4.5 g cm^(-3)

OpenStudy (anonymous):

Divide by 10^3 to go from kg/m^3 to g/cm^3.